OpenStudy (anonymous):

Given the function f(x) = 2(x + 10), find x if f(x) = 24.

OpenStudy (anonymous):

24=2(x+10) 12=x+10 2=x

OpenStudy (anonymous):

sub in 24 instead of f(x) 24 = 2x+20